# uDistil-Whisper: Label-Free Data Filtering for Knowledge Distillation via Large-Scale Pseudo Labelling

### v1

@@ Abstract
Recent work on distilling Whisper’s knowledge into small models using pseudo-labels shows promising performance while reducing the size by up to 50%. This results in small, efficient, and dedicated models. However, a critical step of distillation from pseudo-labels involves filtering high-quality predictions and using only those during training. This step requires ground truth to compare and filter bad examples making the whole process supervised. In addition to that, the distillation process requires a large amount of data thereby limiting the ability to distil models in low-resource settings. To address this challenge, we propose an unsupervised or label-free framework for distillation, thus eliminating the requirement for labeled data altogether. Through experimentation, we show that our best distilled models outperform the teacher model by 5-7 points in terms of WER. Additionally, our models are on par with or better than similar supervised data filtering setup. When we scale the data, our models significantly outperform all zero-shot and supervised models. In this work, we demonstrate that it’s possible to distill large Whisper models into relatively small models without using any labeled data. As a result, our distilled models are 25-50% more compute and memory efficient while maintaining performance equal to or better than the teacher model.
@@ 1 Introduction
In an effort to democratize Automatic Speech Recognition (ASR), significant attention has been given to multilingual models Pratap et al. (2023b); Radford et al. (2023); Communication et al. (2023a); Zhang et al. (2023). These powerful systems, thanks to their massive number of parameters and training data, can simultaneously transcribe in hundreds of languages. That being said, low-resource languages tend to trail behind in performance compared to counterparts like English. For instance, while OpenAI’s latest ASR model Whisper-large-v3 shows a performance in the single digits on English and Spanish, East Asian and African languages show a much lower performance.  [^1] https://github.com/openai/whisper/discussions/1762 While the performance reported on these multilingual models is not as low for Arabic, it is important to highlight that it does not fully and accurately represent the diversity of the different varieties that characterize the Arabic language family Abdul-Mageed et al. (2020); Habash (2022). In addition to the disparity in performance, multilingual models are extremely compute-hungry and are thus not equally accessible to everyone. In order to alleviate this issue, knowledge distillation has been the method of choice for multiple works Hinton et al. (2015); Yang et al. (2023); Sanh et al. (2019); Frantar et al. (2022). It has proven to be powerful, effectively reducing model size (and hence compute and memory requirements) while keeping performance on par with the model originally distilled from. Gandhi et al. (2023a) and Waheed et al. (2024) showcase this for English and Arabic, respectively. That being said, one important limitation of these works is their reliance on a large number of high-quality pseudo labels from the teacher model. Ground-truth labels, which are scarce for low-resource languages, are needed to filter out incorrect teacher predictions and ensure such quality. This constraint thus calls for an unsupervised approach to knowledge distillation.
@@ 3.2.1 Proxy Models
This method utilizes a separate, pre-trained ASR model (referred to as proxy model) to generate reference transcripts for the input speech. The quality of the teacher model’s pseudo-label is then assessed based on the Word Error Rate (pWER) between the proxy model’s reference and the teacher’s reference.
Lower pWER values indicate better agreement between the teacher and proxy models, suggesting higher confidence and reliability in the pseudo-label. A pre-defined WER threshold can be used to filter out examples exceeding the acceptable error rate. In our work, we use SeamlessM4T-large-v2 as our choice of proxy model.
@@ 3.3 Uncertainty Quantification
Entropy. Entropy measures the information uncertainty associated with the teacher’s predicted probability distribution over the vocabulary. High entropy indicates the model is unsure about the correct word, suggesting lower confidence in the pseudo-label. The entropy $H$ for a single utterance with $N$ words is calculated as:
@@
where $c_{i}$ denotes the confidence score assigned by the teacher model for the $i^{\text{th}}$ word.
During filtering, examples with low entropy and a high geometric mean are considered reliable and are retained for training the student model.
@@ 3.3.1 Negative Log Likelihood
We utilize a language model (AceGPT-7B Huang et al. (2024)) to compute the log-likelihood of Whisper output. A lower negative log-likelihood indicates higher compatibility between the pseudo-label and the language model’s understanding of natural language, suggesting a more accurate pseudo-label.
The negative log-likelihood (NLL) for a teacher-predicted pseudo-label sequence (y) with length T is:
@@
where $p(y_{t}|y_{1...t-1})$ is the probability assigned by the language model to the $t-th$ word $(y_{t})$ in the pseudo-label sequence, given the previous words $(y_{1...t-1})$.
Pseudo labels with lower negative log-likelihood values are preferred, as they are more likely to form coherent sequences in the language model, indicating higher quality pseudo labels suitable for training the student models.
@@ 3.3.2 Multimodal Embeddings
In this approach, SONAR Duquenne et al. (2023) is utilized to generate embeddings for both input speech segments and their corresponding pseudo labels. These embeddings capture the contextual information of the speech signals and the text representations derived from the pseudo labels. The similarity between these embeddings is computed using the dot product, serving as a measure of how closely aligned the speech and pseudo-label representations are.
A high similarity score indicates strong contextual alignment between the speech and text, suggesting that the pseudo label accurately reflects the content of the speech segment. This alignment metric is crucial in identifying and retaining high-quality pseudo-labels while filtering out those that do not adequately represent the speech content.
By leveraging the similarity scores derived from SONAR embeddings, we effectively enhance the quality of pseudo labels used in training, as a result improving the overall performance. This method enables more accurate model training in scenarios where labeled data is limited or unavailable, optimizing the efficiency and effectiveness of the learning process.
@@ 3.3.3 Similarity with Synthetic Speech
In this approach, we start by generating synthetic speech for pseudo labels using XTTS-v2  [^2] https://huggingface.co/coqui/XTTS-v2 based on the pseudo label text. Next, we compute the similarity between the synthetic speech and the ground truth speech. To assess the quality of the synthetic speech compared to the original speech segment, we utilize metrics such as PESQ(Perceptual Evaluation of Speech Quality) and MCD (Mel-Cepstral Distortion). A high similarity score suggests that the pseudo labels are likely to have minimal errors.
When evaluating these metrics in relation to the Word Error Rate (WER), we observe that PESQ shows a stronger correlation. Therefore, we select PESQ as the primary metric for our experiments. This choice ensures an effective assessment of synthetic speech quality and alignment with the original speech, enhancing the reliability of our pseudo-labeling approach.
@@ 3.4 Training Data
We follow Waheed et al. (2024) for the data mixture to train our models. As a result, we randomly select segments from a diverse set of datasets, including MGB2 Ali et al. (2019a), MGB3 Ali et al. (2017), FLEURS Conneau et al. (2022), CommonVoice 15.0 Ardila et al. (2020), QASR Mubarak et al. (2021), the Arabic Speech Corpus Halabi (2016), and the Massive Arabic Speech Corpus (MASC) Al-Fetyani et al. (2021). Specifically, we sample 100K and 500K segments, equivalent to approximately 100 and 500 hours of pseudo-labeled speech data, respectively. Our dataset compilation strictly includes only the train splits from each source. We filter out roughly 27% bad quality examples from our training data using the metrics described in Section 3.
@@ 4 Experiments
For fair comparison and analysis, we keep our experimental setup quite identical to Waheed et al. (2024). Apart from evaluating our models on five standard benchmark datasets, and novel dialectal data introduced in Waheed et al. (2024), we extend the evaluation to SADA Alharbi et al. (2024) data. We describe these subsequently.
@@ 4.1 Dataset
FLEURS. The Few-shot Learning Evaluation of Universal Representations of Speech benchmark (FLEURS) Conneau et al. (2022), which is an extension of the FLoRes-101 benchmark Goyal et al. (2022), is an n-way parallel multi-lingual speech dataset comprising 102 languages. Each language is represented by approximately 12 hours of speech. FLEURS is widely utilized for tasks such as automatic speech recognition (ASR), language identification, and speech-to-speech, speech-to-text translation. We use Arabic split which has Modern Standard Arabic (MSA) speech delivered by speakers with Egyptian accents.
@@
Multi-Genre Broadcast. Multi-Genre Broadcast (MGB) Ali et al. (2019a, 2017, b) is an Arabic dataset released under three different editions: MGB2 contains 1,200 hours of speech collected from the TV channel Aljazeera Arabic. It is majoritarily (> 70%) MSA speech and can be categorized as interviews, reports, and conversations. MGB3, on the other hand, is entirely dialectal with 6 hours of Egyptian speech collected from various YouTube channels. The last edition, MGB5, contains 14 hours of speech extracted from YouTube in the Moroccan dialect.
SADA. The Saudi Audio Dataset for Arabic (SADA) Alharbi et al. (2024) is an Arabic dataset comprised of 668 hours of speech (435 of which are labeled) in both spontaneous and read styles. Most of the speech is in different Saudi dialect (Najdi, Hijazi and Khaleeji among others), but it also includes MSA and dialects from other regions, namely the Levant, Yemen, and Egypt. We report the statistics about SADA data in Table 1
@@ 4.2 Baselines
We conduct evaluations on various multilingual and monolingual speech recognition models using different varieties of of Arabic from the datasets mentioned earlier, encompassing both standard and accented MSA, as well as multiple Arabic dialects. We organize these systems into the following categories:
@@ 4.2.2 Zero-Shot Models
Whisper. Whisper Radford et al. (2023) is a versatile multilingual speech model designed for both speech recognition and translation, including Arabic. We assess four Whisper variants: Whisper-Small (W-S), Whisper-Medium (W-M), Whisper-Large v2 (W-L-v2), and Whisper-Large v3 (W-L-v3), using their default decoding parameters with a maximum sequence length of 225 tokens.
SeamlessM4T. SeamlessM4T models excel at generating high-quality transcripts across languages Communication et al. (2023b), but evaluations typically focus on English. To fill this gap, we test three versions of SeamlessM4T—Medium (SM4T-M), Large v1 (SM4T-L-v1), and Large v2 (SM4T-v2)—in a zero-shot setting for Arabic ASR, using the default parameters from the model’s inference pipeline.
@@ 4.2.3 Distilled Models
To compare our results with an equivalent supervised setup, we take two variants (in terms of size) of distilled models presented in Waheed et al. (2024). Apart from reported results, we extend the evaluation on SADA test and valid splits. We take 16-16 and 32-16 models trained with a WER threshold of 80 on 100K and 500K segments.Waheed et al. (2024) find 80 as the optimal threshold which gives the best results. In addition to that we also take models trained without any filtering and treat it as lower bound.
@@ 4.3 Experimental Setup
During evaluations, we adhere to the default decoding parameters specified by the original model implementations unless noted otherwise. Throughout our experiments, we maintain a maximum sequence length of 225 and utilize Word Error Rate (WER) and Character Error Rate (CER) as the primary metrics for assessing performance.
For distillation, we roughly 27% of segments based on the metrics to duplicate the amount of use in supervised setup when $\lambda=80\%$. We refrain from conducting a hyperparameter search due to computational constraints and instead apply the configuration detailed in Gandhi et al. (2023b). The distillation process and its key parameters are detailed in Table 5 in Appendix 1.2.
@@ 5 Results and Discussion
We report the baseline results from Waheed et al. (2024). In addition to that, we evaluate all models on SADA test and validation splits. We do the distillation with two variants of student models. The first, student model has 16 encoder and decoder blocks (UDW-16-16) and the other one has 32 encoder and 16 decoder blocks (UDW-32-16). Following Waheed et al. (2024), we experiment with 100K and 500K segments. We report our main results in 2, and average across different evaluation setups in 3 and top five varieties [^3] We refer dialect labels in SADA as varieties as not all labels are necessarily dialects. of SADA test and validation splits. We report the results on normalized text that includes removing the diacritics. Results on orthographic text can be found in Appendix 1.2.
@@ 1.1 Training Parameters
$warmup\_steps$ $50$
$learning\_rate$ $0.0001$
$lr\_scheduler\_type$ $constant\_with\_warmup$
$batch\_size$ $128$
$max\_label\_length$ $225$
$gradient\_accumulation\_steps$ $1$
@@
Table 5: Training parameters. We use all the default training parameters provided in Huggingface Seq2SeqTrainingArguments unless otherwise specific in this table.

### v2

@@ Abstract
Recent work on distilling Whisper’s knowledge into small models using pseudo-labels shows promising performance while reducing the size by up to 50%. This results in small, efficient, and dedicated models. However, a critical step of distillation from pseudo-labels involves filtering high-quality predictions and using only those during training. This step requires ground truth to compare and filter bad examples making the whole process supervised. In addition to that, the distillation process requires a large amount of data thereby limiting the ability to distil models in low-resource settings. To address this challenge, we propose an unsupervised or label-free framework for distillation, thus eliminating the requirement for labeled data altogether. Through experimentation, we show that our best distilled models outperform the teacher model by 5-7 points in terms of WER. Additionally, our models are on par with or better than similar supervised data filtering setup. When we scale the data, our models significantly outperform all zero-shot and supervised models. We demonstrate that it is possible to distill large Whisper models into relatively small models without using any labeled data. Our distilled models are 25-50% more compute and memory efficient while maintaining performance equal to or better than the teacher model.
@@ 1 Introduction
In an effort to democratize Automatic Speech Recognition (ASR), significant attention has been given to multilingual models Pratap et al. (2023b); Radford et al. (2023); Communication et al. (2023a); Zhang et al. (2023). These powerful systems, thanks to their massive number of parameters and training data, can simultaneously transcribe in hundreds of languages. That being said, low-resource languages tend to trail behind in performance compared to counterparts like English. For instance, while OpenAI’s latest ASR model Whisper-large-v3 shows a performance in the single digits on English and Spanish, East Asian and African languages show a much lower performance [^1] https://github.com/openai/whisper/discussions/1762. While the performance reported on these multilingual models is not as low for Arabic, it is important to highlight that it does not fully and accurately represent the diversity of the different varieties that characterize the Arabic language family Abdul-Mageed et al. (2020); Habash (2022). In addition to the disparity in performance, multilingual models are extremely compute-hungry and are thus not equally accessible to everyone. In order to alleviate this issue, knowledge distillation has been the method of choice for multiple works Hinton et al. (2015); Yang et al. (2023); Sanh et al. (2019); Frantar et al. (2022). It has proven to be powerful, effectively reducing model size (and hence compute and memory requirements) while keeping performance on par with the model originally distilled from. Gandhi et al. (2023a) and Waheed et al. (2024) showcase this for English and Arabic, respectively. That being said, one important limitation of these works is their reliance on a large number of high-quality pseudo labels from the teacher model. Ground-truth labels, which are scarce for low-resource languages, are needed to filter out incorrect teacher predictions and ensure such quality. This constraint thus calls for an unsupervised approach to knowledge distillation.
@@ 3.2.1 Proxy Models
This method utilizes a separate, pre-trained ASR model (referred to as proxy model) to generate reference transcripts for the input speech. The quality of the teacher model’s pseudo-label is then assessed based on the Word Error Rate (pWER) between the proxy model’s reference and the teacher’s reference. Lower pWER values indicate better agreement between the teacher and proxy models, suggesting higher confidence and reliability in the pseudo-label. A pre-defined WER threshold can be used to filter out examples exceeding the acceptable error rate. In our work, we use SeamlessM4T-large-v2 as our choice of proxy model.
@@ 3.3 Uncertainty Quantification
Entropy. Entropy measures the information uncertainty associated with the teacher’s predicted probability distribution over the vocabulary. High entropy indicates the model is unsure about the correct word, suggesting a lower confidence in the pseudo-label. The entropy $H$ for a single utterance with $N$ words is calculated as:
@@
where $c_{i}$ denotes the confidence score assigned by the teacher model for the $i^{\text{th}}$ word. During filtering, examples with low entropy and a high geometric mean are considered reliable and are retained for training the student model.
@@ 3.3.1 Negative Log Likelihood
We make use of an AceGPT-7B Huang et al. (2024) Language Model to compute the log-likelihood of the teacher’s output. A lower negative log-likelihood indicates higher compatibility between the pseudo-label and the language model’s understanding of natural language, suggesting a more accurate pseudo-label. The negative log-likelihood $NLL$ for a teacher-predicted pseudo-label sequence $y$ of length $T$ is:
@@
where $p(y_{t}|y_{1...t-1})$ is the probability assigned by the language model to $y_{t}$, the $t^{\text{th}}$ word in the pseudo-label sequence, given the previous words $(y_{1...t-1})$. Pseudo labels with lower negative log-likelihood values are preferred, as they are more likely to form coherent sequences in the language model, indicating higher quality pseudo labels suitable for training the student models.
@@ 3.3.2 Multimodal Embeddings
In this approach, SONAR Duquenne et al. (2023) is utilized to generate embeddings for both input speech segments and their corresponding pseudo-labels. These embeddings capture the contextual information of the speech signals and the text representations derived from the pseudo-labels. The similarity between these embeddings is computed using the dot product, serving as a measure of how closely aligned the speech and pseudo-label representations are. A high similarity score indicates strong contextual alignment between the speech and text, suggesting that the pseudo-label accurately reflects the content of the speech segment. This helps in identifying and retaining high-quality pseudo-labels while filtering out those that do not adequately represent the speech content. By leveraging the similarity scores derived from SONAR embeddings, we effectively enhance the quality of the pseudo-labels used in training, improving the overall performance as a result. This method enables more accurate model training in scenarios where labeled data is limited or unavailable, optimizing the efficiency and effectiveness of the learning process.
@@ 3.3.3 Similarity with Synthetic Speech
In this approach, we start by generating synthetic speech from the pseudo-labels’ text using XTTS-v2222https://huggingface.co/coqui/XTTS-v2. Next, we compute the similarity between the synthetic speech and the ground truth speech. To assess the quality of the synthetic speech compared to the original speech segment, we utilize metrics such as Perceptual Evaluation of Speech Quality (PESQ) and Mel-Cepstral Distortion (MCD). A high similarity score suggests that the pseudo-labels are likely to have minimal errors. When evaluating these metrics in relation to the Word Error Rate (WER), we observe that PESQ shows a stronger correlation. Therefore, we select PESQ as the primary metric for our experiments. This choice ensures an effective assessment of synthetic speech quality and alignment with the original speech, enhancing the reliability of our pseudo-labeling approach.
@@ 3.4 Training Data
We follow Waheed et al. (2024) for the data mixture to train our models. As a result, we randomly select segments from a diverse set of datasets, including MGB2 Ali et al. (2019a), MGB3 Ali et al. (2017), FLEURS Conneau et al. (2022), CommonVoice 15.0 Ardila et al. (2020), QASR Mubarak et al. (2021), the Arabic Speech Corpus Halabi (2016), and the Massive Arabic Speech Corpus (MASC) Al-Fetyani et al. (2021). Specifically, we sample 100K and 500K segments, equivalent to approximately 100 and 500 hours of pseudo-labeled speech data, respectively. Our dataset compilation strictly includes the train splits only from each source. We filter out roughly 27% bad quality examples from our training data using the metrics described in Section 3.
@@ 4 Experiments
For a fair comparison and analysis, we keep our experimental setup quite identical to Waheed et al. (2024). Apart from evaluating our models on five standard benchmark datasets, and novel dialectal data introduced in Waheed et al. (2024), we extend the evaluation to include the SADA Alharbi et al. (2024) dataset. We describe these corpora below.
@@ 4.1 Dataset
FLEURS. The Few-shot Learning Evaluation of Universal Representations of Speech benchmark (FLEURS) Conneau et al. (2022) is an n-way parallel multi-lingual speech dataset and an extension of the FLoRes-101 benchmark Goyal et al. (2022). It comprises 102 languages, each represented by approximately 12 hours of speech. FLEURS is widely utilized for tasks such as ASR, language identification, and speech-to-speech or speech-to-text translation. We use the Arabic split which has MSA speech delivered by speakers with Egyptian accents.
@@
Multi-Genre Broadcast. Multi-Genre Broadcast (MGB) Ali et al. (2019a, 2017, b) is an Arabic dataset released under three different editions: MGB2 contains 1,200 hours of speech collected from the TV channel Aljazeera Arabic. It is majoritarily MSA speech (> 70%) and is categorized as interviews, reports, and conversations. MGB3, on the other hand, is entirely dialectal with 6 hours of Egyptian speech collected from various YouTube channels. The last edition, MGB5, contains 14 hours of speech extracted from YouTube in the Moroccan dialect.
SADA. The Saudi Audio Dataset for Arabic (SADA) Alharbi et al. (2024) is an Arabic dataset comprised of 668 hours of speech (435 of which are labeled) in both spontaneous and read styles. Most of the speech is in different Saudi dialects (Najdi, Hijazi and Khaleeji among others), but it also includes MSA and dialects from other regions, namely the Levant, Yemen, and Egypt. We report the statistics about the SADA data in Table 1
@@ 4.2 Baselines
We conduct evaluations on various multilingual and monolingual speech recognition models using different varieties of Arabic from the datasets mentioned earlier, encompassing both standard and accented MSA, as well as multiple Arabic dialects. We organize these systems into the following categories:
@@ 4.2.2 Zero-Shot Models
Whisper. Whisper Radford et al. (2023) is a versatile multilingual speech model designed for both speech recognition and translation, including Arabic. We assess four Whisper variants: whisper-small (W-S), whisper-medium (W-M), whisper-large-v2 (W-L-v2), and whisper-large-v3 (W-L-v3), using their default decoding parameters with a maximum sequence length of 225 tokens.
SeamlessM4T. SeamlessM4T models excel at generating high-quality transcripts across languages Communication et al. (2023b), but evaluations on them typically focus on English. To fill this gap, we test three of its versions – seamless-m4t-medium (SM4T-M), seamless-m4t-v1-large (SM4T-L-v1), and seamless-m4t-v2-large (SM4T-L-v2) – under a zero-shot setting for Arabic ASR, using the default parameters from the model’s inference pipeline.
@@ 4.2.3 Distilled Models
To compare our results with an equivalent supervised setup, we take two variants (in terms of size) of distilled models presented in Waheed et al. (2024). Apart from the reported results, we extend the evaluation on SADA’s test and valid splits. We take 16-16 and 32-16 models trained with a WER threshold of 80 on each of 100K and 500K segments. Waheed et al. (2024) find 80 as the optimal threshold which gives the best results. We also take models trained without any filtering and treat those as a lower bound.
@@ 4.3 Experimental Setup
During evaluations, we adhere to the default decoding parameters specified by the original model implementations unless noted otherwise. Throughout our experiments, we maintain a maximum sequence length of 225 and use WER and CER as the primary metrics for assessing performance.
For distillation, we select roughly 27% of the top segments based on the metrics, to duplicate the amount used in a supervised setup when $\lambda=80\%$. We refrain from conducting hyperparameter search due to computational constraints, and instead apply the configuration detailed in Gandhi et al. (2023b). The distillation process and its key parameters are detailed in Table 5 in Appendix 1.2.
@@ 5 Results and Discussion
We report the baseline results from Waheed et al. (2024) and evaluate all models on SADA’s test and validation splits. We do the distillation with two variants of student models. The first student has 16 of each of the encoder and decoder blocks (UDW-16-16) and the second one has 32 encoder and 16 decoder blocks (UDW-32-16). Following Waheed et al. (2024), we experiment with 100K and 500K segments. We report our main results in Table 2, and the averages across different evaluation setups in Table 3. Results on the top five varieties [^3] We refer dialect labels in SADA as varieties as not all labels are necessarily dialects. of SADA are provided in Table 4. The reported results are computed on normalized text that includes removing diacritics. Results on the orthographic text (no normalization) can be found in Appendix 1.2.
@@ 1.1 Training Parameters
warmup_steps $50$
learning_rate $0.0001$
lr_scheduler_type constant_with_warmup
batch_size $128$
max_label_length $225$
gradient_accumulation_steps $1$
@@
Table 5: Training parameters. All the training parameters are the default ones provided in Huggingface Seq2SeqTrainingArguments unless specified otherwise in this table.